# Non-perturbative self-consistent electron-phonon spectral functions and transport

### v1

@@ Abstract
The spectral function is a key quantity in many-body physics, linking theory and experiment. Experimentally, spectral functions can be extracted using high-resolution angle-resolved photoemission spectroscopy [1]. Theoretically, spectral functions can be calculated from Green’s functions, including many-body correlation via the self-energy. The coupling between electrons and phonons, the collective vibration of the lattice, plays a crucial and often dominant role in shaping the electron spectral function [2]. The electron-phonon interaction leads to the broadening and renormalization of the quasiparticle peak [3, 4, 5, 6, 7], photoemission kinks in metals [8, 9], and additional structures such as satellite peaks [10, 11, 12, 13, 14, 15]. Understanding and predicting these effects is the foundation of electron-phonon physics.
@@
Ab initio calculation of the electron-phonon self-energy and spectral functions is usually limited to perturbation theory. Standard approaches use the perturbative Fan–Migdal (FM) self-energy [16, 17] and solve the Dyson equation [10] or apply the cumulant expansion [18, 19, 20, 21, 22, 15, 23, 24, 25] to calculate the spectral functions. However, perturbative approaches often fail, leading to divergent quasiparticle dispersion [26, 27, 25] or unphysical spectral properties [25]. Importantly, this problem is not restricted to strongly coupled systems. The negative curvature in the quasiparticle dispersion of the Fröhlich model [26] and the divergent linewidth of piezoelectric materials [28] are examples where perturbation theories fail. More sophisticated methods that include higher-order electron-phonon coupling [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43] are limited to model Hamiltonians. Non-perturbative methods based on molecular dynamics [44, 45] or special-displacement supercells [46, 47] are difficult to converge [45, 48], and cannot describe non-adiabatic electron-phonon coupling which is crucial in polar materials [27, 49].
@@
The electron spectral function also describes the transport of charged carriers. Using the Green–Kubo (GK) theory [50, 51], one can compute the conductivity without current vertex correction directly from the spectral functions [52, 53], which is sometimes called the Kubo formula in the bubble approximation or the independent-particle approximation [54]. This formulation has the advantage of including broadening and satellites of spectral functions beyond the quasiparticle approximation [55, 56, 57, 58, 45].
@@
A material of particular interest for the non-perturbative electron-phonon coupling is monolayer InSe. While monolayer InSe is known for its high electron mobility [59, 60], its valence band has a flat, Mexican-hat-like dispersion [61, 62, 63, 64, 65, 66, 67], which leads to strong electron-phonon interaction. In particular, state-of-the-art one-shot calculations of the spectral function yield a dynamical splitting of a single valence band into two bands with a gap larger than 150 meV [68]. The hole mobility of monolayer InSe has been studied using the Boltzmann transport equation (BTE) [69, 70, 71] and the GK method with the one-shot FM spectral function [68]. Since the renormalization is large, non-perturbative corrections are expected to play an important role in InSe.
@@
In this work, we present a non-perturbative ab initio calculation of the electron-phonon spectral function and conductivity. We use the self-consistent Fan–Migdal (scFM) method [35, 36, 37, 38, 39, 40, 41, 42, 43], which self-consistently updates the self-energy with the renormalized Green’s function. Using the self-consistent spectral functions, we then use the GK method to compute the conductivity. We find that self-consistency significantly affects the spectral function, direct current (DC) conductivity, and optical conductivity of monolayer InSe. Such an effect is present in any materials and can be studied using this scFM method.
The most used approximation for the electron-phonon self-energy is the FM formula [2]:
@@
where is the volume of the Brillouin zone, the electron-phonon matrix element, and the electron and phonon energies, the Fermi–Diarc distribution function, , and the Bose-Einstein distribution function. We neglect the band-off-diagonal self-energy, which is negligible if the bands do not hybridize [74], as studied here. In this work, we assume that the bare electron Green’s function is obtained from Kohn–Sham density functional theory and that the phonon Green’s function and electron-phonon coupling are obtained from adiabatic density functional perturbation theory [75, 76, 2]. We neglect the renormalization of the phonon Green’s function.
@@
Using the FM self-energy, the renormalized electron dispersion can be computed at the level of Rayleigh–Schrödinger (RS) perturbation theory [77, 78, 27, 2, 23] by adding the on-shell FM self-energy to the bare band energy:
@@
The energy-dependent spectral function can be calculated by solving the Dyson equation:
@@
To capture higher-order electron-phonon coupling, one often uses the cumulant approximation, which replaces the Dyson equation with the cumulant expansion
@@
where is the Heaviside function. One then approximates the cumulant function using second-order perturbation theory [21]:
@@
Unlike the FM approximation, which gives a single satellite [15], the cumulant approximation can generate multiple satellite peaks as observed in experiments [11, 12, 13, 14]. However, the cumulant approximation can lead to unphysically large renormalizations for -points away from the band edge [25, 43], as does the RS [26] approximation.
@@
To include the non-perturbative effects of electron-phonon coupling, we go one step further and use the self-consistent Fan–Migdal (scFM) method. The scFM self-energy is obtained by replacing the bare electron Green’s function in the FM approximation with the dressed one. The imaginary part of the scFM self-energy is [57]:
@@
The real part is then obtained from the Kramers–Kronig relation:
@@
where denotes the Cauchy principal value. The spectral function in Eq. (7) is obtained by solving the Dyson equation:
@@
A self-consistent solution of Eqs. (7-10) gives the scFM self-energy and spectral function. For metals or doped semiconductors, we iteratively update the temperature-dependent chemical potential to match the total electron occupation.
@@
We remark that it is advantageous to calculate the imaginary part first and then compute the real part using the Kramers–Kronig relation. In fact, the relation between the real parts of and involves a -dependent energy integral, see Eq. (S46) of the supplemental of Ref. [28]. The separation of Eq. (7) and Eq. (8) allows to perform the -integration for the imaginary part first, and then perform a -independent energy integral after.
@@
In practical ab initio calculations, several hundred bands per atom are needed to converge the real part of the FM self-energy [48]. Calculating such a term via the Kramers–Kronig transformation is impractical since it requires calculating up to eV. We simplify the calculation by exploiting the fact that the contribution of high-energy bands to the self-energy at small can be viewed as a static correction, which can be computed via the self-consistent solution of linear Sternheimer equations [77, 78]. We self-consistently compute considering only the states inside the active energy window . We approximate the remainder to be static and compute it using Wannier function perturbation theory (WFPT) [79]:
Here, is the FM self-energy [Eq. (1)] from states outside the active space, , evaluated at the bare eigenvalue , and is the Debye–Waller self-energy [5, 4, 3].
@@
In practice, we initialize the scFM self-energy with from Eq. (1) with meV. We iterate until the maximum change of the self-energy for all states and energies (with respect to the band edge) is below meV. We have implemented scFM as an extension of our ElectronPhonon.jl package [80], which relies on quantities computed using Quantum ESPRESSO [81], Wannier90 [82], and EPW [83, 84]. Computational details are provided in Sec. S1 of the SM [72].
@@
The scFM method, also known as the self-consistent Born approximation or the noncrossing approximation, has been applied to the Holstein [35, 36, 37, 39, 40, 41, 42, 43] and Anderson–Holstein models [38]. State-of-the-art ab initio calculations [57, 58] used a one-shot calculation of Eqs. (7) and (8) to perturbatively compute the electron-phonon self-energy using the spectral function obtained from dynamical mean-field theory calculations. Here, we self-consistently solve Eqs. (7-10) to include non-perturbative effects of electron-phonon coupling on the spectral function in real materials.
@@
We apply the scFM method to monolayer InSe which has a Mexican-hat-like valance band [61, 62, 63, 65]. Due to this dispersion, hole-doped InSe monolayers exhibit strong electron-phonon effects. In particular, the one-shot FM spectral functions show a dynamic splitting of a single valence band into two peaks with a gap of 150 meV or larger [68]. We study whether this large renormalization persists under a self-consistent update of the spectral function.
@@
Figures 1(a-d) show the valence band spectral functions of monolayer InSe at 100 K which is compared to the monolayer angle-resolved photoemission spectra (ARPES) intensity from Ref. [73] measured at the same temperature. We first find that the RS dispersion has an inverted curvature at the VBM, an artifact also observed for the Fröhlich model with strong electron-phonon coupling [26]. The cumulant spectral function shows a single broad peak with no satellites and a prominent dip at . The one-shot FM method yields two dynamically split bands with a splitting of 200 meV at the VBM, consistent with Ref. [68]. However, this large splitting vanishes under self-consistency. The scFM spectral function shows a broadened quasiparticle peak with multiple satellites. The distance between the satellite peaks ranges from 30 to 50 meV and is comparable to the LO phonon energy of 27 meV. An in-depth analysis of the difference between FM and scFM for different temperatures is provided in Sec. S2 of the SM [72]. Our scFM results are consistent with the absence of a 150 meV scale gap [65, 73] and the flatness of the band shown in Fig.1(e), in better agreement than the cumulant result. We expect that higher-resolution ARPES will observe the satellites predicted by our scFM calculation.
@@
Beyond ARPES predictions that can sometimes be challenging to verify experimentally, the electron spectral function can be directly used to compute carrier conductivity, which is a much simpler quantity to measure. Using the GK linear response theory and neglecting the current vertex correction, the conductivity is [52]
@@
The corresponding mobility is with the carrier concentration. Equation (12) has been used to calculate the conductivity due to electron-electron interaction [53], electron-phonon interaction [55, 56, 45], or both [57, 58]. Here, we study how the self-consistent renormalization of the spectral function affects the DC and optical conductivity.
@@
Figure 2(a) shows the intrinsic DC mobility of InSe. While the Boltzmann transport equation (BTE) and self-energy relaxation time approximation (SERTA) mobilities [85] are almost the same and agree with those of Ref. [70, 71], the GK mobility with the scFM spectral function is 3–4 times larger. This result implies that the quasiparticle approximation leads to a significant underestimation of the mobility. We also find that using the one-shot FM spectral function instead of the scFM leads to an overestimation by more than an order of magnitude. These results can be understood in terms of the effective electron lifetimes shown in Fig. 2(b). The SERTA and BTE mobilities are linked to the inverse electron lifetimes, which are proportional to the imaginary part of the FM self-energy at the bare band energy,
For the GK mobility, we define the effective inverse lifetimes using the self-energy at the quasiparticle energy instead of the bare band energy:
@@
Here, the quasiparticle energy and are the energy at which the corresponding spectral functions [Eq. (3)] and [Eq. (9)] are maximal. Near the VBM, the BTE inverse lifetimes have a peak, while the scFM ones show a smoother variation, due to self-consistent broadening. The lower scattering rates increase the mobility. For the one-shot FM self-energy, the scattering rates near the VBM are tiny, because the renormalized valence bands lie within the DFT band gap where the one-shot FM self-energy has a negligible imaginary part. This is an artifact of the one-shot calculation.
@@
Finally, we show in Fig. 3 the alternating current (AC) mobility using the FM and scFM self-energy and the GK method. For the one-shot FM spectral function, we find a peak around 200–300 meV, corresponding to the gap shown in Fig. 1(c) [68]. This feature is found to be stable regardless of the broadening parameter (see Fig. LABEL:supp-fig:ac_sc_broadening of the SM [72]). However, when the scFM spectral function is used, the unphysical peaks are washed out due to the reduction of the splitting energy and the broadening of the Drude peak.
@@
The self-consistent framework developed in this work provides a versatile tool for analyzing the effect of higher-order electron-phonon coupling with moderate computational effort, filling the gap between single-shot perturbative approximations and numerically exact methods. While we have focused on the spectral functions and conductivity, our framework can be extended to other properties such as superconductivity and indirect absorption [86, 87]. Our work lays the groundwork for applying diagrammatic many-body approaches to electron-phonon systems. The scFM method could be extended to study contributions beyond the Fan–Migdal diagram, such as the dynamical screening of phonons [35, 36, 88, 89, 90, 91] and the electron-phonon vertex correction [39, 41, 92] in real materials. Combining the scFM formalism with dynamical mean-field theory [93, 94] and its diagrammatic extensions [95] to study the interplay of electron-electron [53] and electron-phonon coupling [96] on transport beyond the one-shot approximation [57, 58] is another promising direction for future research.

### v2

@@ Abstract
Spectral functions are essential in many-body physics, linking theory and experiment. Experimentally, spectral functions can be extracted using high-resolution angle-resolved photoemission spectroscopy [1]. Theoretically, spectral functions can be calculated from Green’s functions, including many-body correlation via the self-energy. The coupling between electrons and phonons, the collective vibration of the lattice, plays a crucial and often dominant role in shaping the electron spectral function [2]. The electron-phonon (-ph) interaction leads to the broadening and renormalization of the quasiparticle peak [3, 4, 5, 6, 7], photoemission kinks in metals [8, 9], and additional structures such as satellite peaks [10, 11, 12, 13, 14, 15]. Understanding and predicting these effects is the foundation of electron-phonon physics.
@@
Ab initio calculation of the -ph self-energy and spectral functions is usually limited to perturbation theory. Standard approaches use the perturbative Fan–Migdal (FM) self-energy [16, 17], solving the Dyson equation [10] or applying the cumulant expansion [18, 19, 20, 21, 22, 15, 23, 24, 25, 26] for spectral functions. However, these approaches often fail, causing divergent quasiparticle dispersion [27, 28, 26] or unphysical spectral properties [26]. Notably, this problem is not limited to strongly coupled systems: the negative curvature in the quasiparticle dispersion of the Fröhlich model [27] and the divergent linewidth of piezoelectric materials [29] illustrate failures at all coupling strengths. Advanced methods incorporating higher-order -ph interactions [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44] are limited to model Hamiltonians. Non-perturbative methods using molecular dynamics [45, 46] or special-displacement supercells [47, 48] are difficult to converge [46, 49], and cannot capture non-adiabatic -ph coupling, crucial in polar materials [28, 50].
@@
The electron spectral function also governs charge transport. Within the Green–Kubo theory [51, 52], neglecting current vertex corrections allows direct computation of conductivity from spectral functions [53, 54]. This approach, known as the bubble approximation, has the advantage of including broadening and satellites beyond the quasiparticle approximation [24, 55, 56, 57, 46].
@@
A material of particular interest for non-perturbative -ph coupling is monolayer InSe. While known for high electron mobility [58, 59], its valence band features a flat, Mexican-hat-like dispersion [60, 61, 62, 63, 64, 65, 66], leading to strong -ph interaction. In particular, state-of-the-art one-shot calculations predicted a dynamical splitting of the valence band into two peaks in the spectral function with a gap larger than 150 meV [67]. The hole mobility of monolayer InSe has been studied using the Boltzmann transport equation (BTE) [68, 69, 70] and the bubble approximation with the one-shot spectral function [67]. Yet, given the strong renormalization, non-perturbative corrections are expected to be essential in monolayer InSe.
@@
In this work, we present a non-perturbative ab initio calculation of the -ph spectral function and conductivity. Our approach, the self-consistent (sc) method [36, 37, 38, 39, 40, 41, 42, 43, 44], extends the FM self-energy to include the self-consistently renormalized spectral functions, analogously to the sc method for electron-electron interactions [71]. Using self-consistent spectral functions, we compute conductivity using the bubble approximation. We find that self-consistency significantly affects the spectral function and conductivity of monolayer InSe. This effect is present in any materials and can be explored with the sc method.
The most common approximation for the -ph self-energy is the FM formula [2]:
@@
where is the Brillouin zone volume, the -ph matrix element, and the electron and phonon energies, the Fermi–Dirac distribution, , and the Bose-Einstein distribution. We neglect the band-off-diagonal self-energy, which is a valid approximation if the bands do not hybridize [74], as in InSe. We refer to Eq. (1) as the -FM self-energy, corresponding to a one-loop diagram (see Fig. LABEL:supp-fig:feynman(a) of the SM [72]) with bare electron and phonon Green’s functions, and , as in the approximation [75]. In this work, we assume that is obtained from Kohn–Sham density functional theory and that and -ph coupling come from adiabatic density functional perturbation theory [76, 77, 2]. We neglect the renormalization of the phonon Green’s function since we focus on the intrinsic and low doping regime where the phonon renormalization is small [2]. The total self-energy also contains the Debye–Waller (DW) term [5, 4, 3]:
In this work, we compute the DW term within the rigid-ion approximation [78].
@@
Using the self-energy, the renormalized electron dispersion can be computed within Rayleigh–Schrödinger (RS) perturbation theory [79, 80, 28, 2, 23] as
@@
The energy-dependent spectral function is computed by solving the Dyson equation:
with the chemical potential.
@@
To capture higher-order -ph coupling, one often uses the cumulant approximation. It replaces the Dyson equation with the expansion
@@
where is the Heaviside function and the cumulant function. We adopt the retarded cumulant method [81, 82, 24] as we deal with finite temperatures. The cumulant function is approximated via second-order perturbation theory [81, 21, 82]:
@@
Unlike the approximation, which only produces a single satellite [15], the cumulant approach can generate multiple satellite peaks, as seen in experiments [11, 12, 13, 14]. However, it has been shown that the cumulant approximation can lead to unphysically large renormalizations for points away from the band edge [26, 44], similar to the RS approximation [27].
@@
To incorporate non-perturbative -ph effects, we adopt the sc method, where the self-energy is obtained by replacing the bare electron Green’s function in the approximation with the dressed one (see Fig. LABEL:supp-fig:feynman(b) of the SM [72]). The imaginary part of the sc self-energy is given by [56]
@@
The real part is obtained using the Kramers–Kronig relation, with the addition of the real-valued DW term:
@@
where denotes the principal value. The spectral function in Eq. (8) is obtained via the Dyson equation:
@@
Solving Eqs. (8)-(11) self-consistently yields the sc self-energy and spectral function. We also iteratively adjust the temperature-dependent chemical potential to match the total electron occupation.
@@
We remark that it is advantageous to first compute the imaginary part and then use the Kramers–Kronig relation to obtain the real part. The relation between the real parts of and involves a -dependent energy integral [56]. Separating Eq. (8) and Eq. (9) enables performing the -integration for the imaginary part first, followed by a -independent energy integral.
@@
In practical ab initio calculations, several hundred bands per atom are needed to converge the real part of the FM self-energy [49]. Calculating this via the Kramers–Kronig transformation is impractical, as it demands evaluating up to eV. We simplify the calculation by treating the contribution of high-energy bands as a static correction. We thus self-consistently compute only for the states inside the active space window , and add the static correction as
Here, is the one-shot FM self-energy [Eq. (1)] from states outside the active space, , evaluated at the bare eigenvalue . This term can be efficiently computed by solving the linear Sternheimer equation [79, 80] and interpolating using Wannier function perturbation theory (WFPT) [83]. The static corrections in the second line of Eq. (12) are always included during the self-consistent calculation of the sc self-energy, i.e., Eq. (12) replaces Eq. (9). (See Sec. LABEL:supp-sec:active_space of the SM [72] for additional details.)
@@
In practice, we initialize the sc self-energy as a constant with meV. We do not use any artificial broadening during the sc iteration, and the choice of does not affect the converged sc self-energy. We iterate until the maximum change of is below meV. The sc method is implemented as an extension of our ElectronPhonon.jl package [84], which uses quantities computed from Quantum ESPRESSO [85], Wannier90 [86], and EPW [87, 88]. Computational details are provided in Sec. LABEL:supp-sec:comp_details of the SM [72].
@@
The sc method, also known as the self-consistent Migdal, self-consistent first-Born, or the non-crossing approximation, has been applied to the Holstein [89, 90, 36, 37, 38, 40, 41, 42, 43, 44] and Anderson–Holstein models [39]. This non-perturbative method sums the infinite series of non-crossing rainbow diagrams (see Fig. LABEL:supp-fig:feynman(c) of the SM [72]), including those with an arbitrary number of -ph vertices. It is expected to be valid from weak to intermediate coupling but will fail in the strong coupling regime, as in the case of self-trapped small polarons [91]. State-of-the-art ab initio calculations [56, 57] used a one-shot calculation of Eqs. (8, 9) (without ) with the spectral function obtained from dynamical mean-field theory to compute the -ph self-energy. Here, we self-consistently solve Eqs. (8-12) to include non-perturbative -ph effects on the spectral function.
@@
We apply the sc method to monolayer InSe, which has a Mexican-hat-like valance band [60, 61, 62, 64]. This dispersion leads to strong -ph effects in hole-doped InSe monolayers: the one-shot spectral functions show a dynamic splitting of 150 meV or larger [67]. We study whether this large renormalization persists under a self-consistent update of the spectral function.
@@
Figures 1(a-d) show the valence band spectral functions of monolayer InSe at 100 K, compared to the angle-resolved photoemission spectroscopy (ARPES) photointensity [73] measured at the same temperature. The RS dispersion has several kinks with inverted curvature near the valence band maximum (VBM), an artifact also observed for the Fröhlich model [27]. The cumulant spectral function shows a single broad peak with no satellites and a prominent dip at . The one-shot method yields two dynamically split bands with a splitting of 170 meV, consistent with Ref. [67]. However, this large splitting vanishes under self-consistency. The sc spectral function shows a broadened quasiparticle peak with multiple satellites, with a distance between the peaks of 30–50 meV, comparable to the LO phonon energy of 27 meV. An analysis of the difference between and sc self-energies is provided in the End Matter (Fig. 4). Our sc result is consistent with the absence of a 150 meV scale gap in the ARPES photointensity [64, 73]. Although both the sc and cumulant spectral functions are consistent with the experimental observation within the experimental resolution, higher-resolution ARPES may observe the satellites predicted by our sc calculation.
@@
Beyond ARPES predictions that can sometimes be challenging to verify experimentally, the electron spectral function can be directly used to compute carrier conductivity, a simpler quantity to measure. Neglecting the current vertex corrections, we calculate the conductivity in the bubble approximation, which is given by [53, 54]
@@
The corresponding mobility is with the carrier concentration. Equation (13) has been used to calculate the conductivity due to electron-electron interaction [54], -ph interaction [24, 55, 46], or both [56, 57]. Here, we investigate how the self-consistent renormalization of the spectral function affects the direct current (DC) and alternating current (AC) mobility of monolayer InSe.
To study transport in the intrinsic limit, we set a small carrier concentration, placing the chemical potential within the band gap. In this nondegenerate regime, mobility should be independent of carrier concentration [92, 93]. However, as detailed in the End Matter, we find that both the and cumulant spectral functions yield strongly -dependent mobility (Fig. 5). We attribute this artifact to the slow decay of the spectral function above the band edge, which induces a large number of carriers on the tail of the spectral function. To address this problem, we developed a truncation scheme that removes the tail above a cutoff energy (see Sec. LABEL:supp-sec:decayof the SM [72]), which results in a -independent mobility. However, this scheme cannot be applied to the cumulant spectral function of InSe due to its large broadening. The sc method does not suffer from this problem and yields a -independent mobility.
@@
Figure 2(a) shows the intrinsic DC mobility of InSe. While the Boltzmann transport equation (BTE) and self-energy relaxation time approximation (SERTA) mobilities [94] are nearly identical and agree with the results from Refs. [69, 70], the bubble mobility with the sc spectral function is 3–4 times larger. This suggests that the quasiparticle approximation significantly underestimates the mobility. Using the one-shot spectral function instead of the sc one leads to an overestimation by more than an order of magnitude. These results can be understood in terms of the effective electron scattering rates shown in Fig. 2(b), which are inversely proportional to the mobility. The SERTA and BTE mobilities are linked to the the imaginary part of the self-energy at the bare band energy:
For the spectral functions, we define the effective scattering rates using the self-energy at the renormalized quasiparticle energy:
@@
The renormalized quasiparticle energies, and , are the energies where the spectral functions and are maximal. Near the VBM, the BTE scattering rates exhibit a peak, while the sc result shows a smoother variation and smaller values due to the self-consistent broadening. For the one-shot self-energy, the scattering rates near the VBM are tiny because the renormalized valence bands lie within the DFT band gap where the imaginary part of the self-energy is small. This behavior is an artifact of the one-shot approximation.
@@
Finally, we show in Fig. 3 the AC mobility calculated using the and sc spectral functions. For the one-shot spectral function, we observe a peak around 150–250 meV, corresponding to the gap shown in Fig. 1(c), as in Ref. [67]. This feature persists with a larger artificial broadening (see Fig. LABEL:supp-fig:ac_sc_broadening of the SM [72]). However, with the self-consistent spectral function, the unphysical peaks are smoothed out, due to the reduced splitting and the broader Drude peak.
@@
The self-consistent framework developed in this work offers a versatile tool for analyzing the non-perturbative effect of -ph coupling with moderate computational effort, bridging the gap between one-shot approximations and numerically exact methods. While we have focused on the spectral functions and conductivity, this framework can be extended to other properties such as superconductivity and indirect absorption [95, 96]. Our work also lays the groundwork for applying diagrammatic many-body approaches to -ph systems. The sc method could be extended to explore contributions beyond the Fan–Migdal diagram. Notably, the development of a fully self-consistent scheme using the renormalized phonon Green’s function [36, 37, 97], the inclusion of the dynamical screening [98, 99, 100], and the consideration of -ph vertex correction [40, 42, 101] are promising avenues for future work. Combining the sc formalism with dynamical mean-field theory [102, 103] and its diagrammatic extensions [104] to study transport from the interplay of electron-electron [54] and -ph coupling [105] beyond the one-shot approximation [56, 57] is another interesting direction for future research.
The source code and data associated with this work are available on the Materials Cloud Archive [106].